scores on acommon final examination in alarge enrollment, multiple selection freshman course are normally distributed with a mean of 72. 7 and standard deviation of 13.1 a. Find the probability that the score X on a randomkly selected examination paper is between 70 and 80


1
Expert's answer
2022-04-06T09:30:07-0400

We have a normal distribution, "\mu=72.7, \sigma=13.1."

Let's convert it to the standard normal distribution, "z=\cfrac{x-\mu}{\sigma};"

"z_1=\cfrac{70-72.7}{13.1}=-0.21, \\z_2=\cfrac{80-72.7}{13.1}=0.56,"


"P(70<X<80)=P(-0.21<Z<0.56)=\\=P(Z<0.56)-P(Z<-0.21)="

"=0.71226-0.41683=0.29543" ​(from z-table).